Four forces act on an object, given by A = 4 N due East, B = 5 N due North, C = 7 N due West, and D = 9 N due South. What is the magnitude and direction of the net force on the object?

A) 25 N due South

B) 7 N due South

C) 5 N at 53.14 degrees South of West

D) 5 N at 36.86 degrees South of West

E) None of the above
Answer #1

The net force along x axis

Fx = A- C= 4 N - 7 N = - 3 N

along y axis

Fy = B- D = 5N - 9 N = - 4 N

magnitude of force

F = sqrt Fx^2 + Fy^2

= sqrt ( -3)^2 + (-4)^2 = 5 N

theta = tan^-1 ( Fy/ Fx ) = tan^-1 ( -4/-3) = 53.14 degrees south of west

option (C) is correct answer